a survey unofficially claimed that in every five (5) young executives , only one (1) practices good reading habits.

what is the probability that out of 10 10 young executives , two executives practice good reading habits


1
Expert's answer
2022-03-20T14:51:17-0400

The pribability that aexecutive reads habits good:

"p = \\frac{1}{5} \\Rightarrow q = 1 - p = \\frac{4}{5}p= \n5\n1\n\u200b\n \u21d2q=1\u2212p= \n5\n4"

Then, according to the Bernoulli formula, the wanted probability is

"P_{10}(2)=C _{10}^2p^2 q^8
 =\frac{10!}{ 2!8!} (\frac{1}{5})^2(\frac{4}{5})^8=\frac{580824}{1953155}=0.301989888"